# Agreement-Based Cascading for Efficient Inference

### v3

@@ Abstract
Adaptive inference schemes reduce the cost of machine learning inference by assigning smaller models to easier examples, attempting to avoid invocation of larger models when possible. In this work we explore a simple, effective adaptive inference technique we term Agreement-Based Cascading (ABC). ABC builds a cascade of models of increasing size/complexity and uses agreement between ensembles of models at each level of the cascade as a basis for data-dependent routing. Although ensemble execution introduces additional expense, we show that these costs can be easily offset in practice due to large expected differences in model sizes, parallel inference execution capabilities, and accuracy benefits of ensembling. We examine ABC theoretically and empirically in terms of these parameters, showing that the approach can reliably act as a drop-in replacement for existing models and surpass the best single model it aims to replace in terms of both efficiency and accuracy. Additionally, we explore the performance of ABC relative to existing cascading methods in three common scenarios: (1) edge-to-cloud inference, where ABC reduces communication costs by up to 14×14$\times$14 ×; (2) cloud-based model serving, where it achieves a 3×3$\times$3 × reduction in rental costs; and (3) inference via model API services, where ABC achieves a 222-25×25$\times$25 × reduction in average price per token/request relative to state-of-the-art LLM cascades.
@@ 1 Introduction
  • We empirically evaluate ABC on a wide range of image and language tasks and find that ABC not only improves efficiency, but also accuracy, compared to the model that it aims to replace. We then consider the performance of ABC relative to existing cascading methods in common inference scenarios, including (1) edge-to-cloud inference where ABC reduces communication costs by up to 14×14$\times$14 ×, (2) model-serving on heterogeneous GPUs, where ABC reduces rental costs by up to 3×3$\times$3 × and (3) inference using black-box access to model API services, where ABC shows up to a 25×25$\times$25 × reduction in average price per token.
@@ 4.2 Deferral Rule-Based Drop-in Cascade
A deferral-based cascade consists of a finite set of models and a deferral rule. Here, the idea is to start with the most resource-efficient model and use the deferral rule to determine if a better-performing model with a higher resource cost should be used for the current sample. The deferral rules themselves are designed to have negligible inference cost given the models inference outputs. For this exposition, we will restrict ourselves to cascades with only two levels, though the discussion in this section can be readily generalized to larger cascades. Let the cascade consist of an ensemble $H^{k}_{1}$ at the lower level, and the larger model $h_{2}$ at the higher level; =$\mathcal{M}={\{H^{k}_{1},h_{2}\}}$. Let $r(x)$ denote a deferral rule;
@@
In this work, we focus on deployment scenarios where a decrease in model performance is a critical concern. This is particularly relevant when cascades are intended to function as a drop-in replacement for an existing deployment of an expensive model. This notion can be formulated as maximizing the number of calls to the smaller model while retaining the accuracy of the larger model. With a slight abuse of notation, let $\mathcal{M}_{r}(x)$ denote the prediction of the cascade with =$\mathcal{M}={\{H^{k}_{1},h_{2}\}}$ and deferral rule $r$. Then we desire that for a choice of a small error budget $\xi>0$,
@@ 4.3 Deferral using Ensemble Agreement
Assuming access to a safe deferral rule, observe that for $\xi\geq\epsilon>0$, such rules are feasible for the program in Equation 2 (see Proposition 4.1). This implies that every cascade, =$\mathcal{M}={\{H_{1}^{k},h_{2}\}}$ using a safe deferral rule $r$, is competitive with the larger model $h_{2}$ in terms of accuracy. We empirically evaluate these drop-in cascades in §5. Note that the optimal safe deferral rule can lead to an improvement in the accuracy of ABC in theory (see, Appendix A), and we see accuracy improvements in our experiments (§5).
@@
$r_{\text{v}}(x;\theta_{v})$ =$=\begin{cases}1&\text{vote}(x;H^{k}_{1})\leq\theta_{v}\\ 0&\text{otherwise}.\end{cases}$ (3)
$r_{\text{s}}(x;\theta_{s})$ =$=\begin{cases}1&s(x;H^{k}_{1})\leq\theta_{s}\\ 0&\text{otherwise}.\end{cases}$ (4)
@@ 4.4 Inference Cost Savings and Competitiveness
Let =$\mathcal{M}={\{H_{1}^{k},h_{2}\}}$ be two classifiers and $r$ a deferral rule such that $r$ is a safe deferral rule for $H_{1}^{k}$ according to Definition 4.1, for a distribution $\mathbb{P}$ over $\mathcal{X}\times\mathcal{Y}$. Then for every $\xi\geq\epsilon>0$, the agreement based cascading (ABC) classifier $\mathcal{M}_{r}(x)$ is such that,
@@
[…]
@@ 5 Experiments
This section evaluates our training-free Agreement-Based Cascading (ABC) approach across a variety of language and vision tasks, focusing on accuracy and inference cost. First, in §5.1.1, we examine ABC’s accuracy-cost tradeoff against state-of-the-art models under full parallelization, setting the parallelization factor $(\rho)$ to $\rho=1.0$ for optimal tradeoff. Next, in §5.1.2, we show that using ensembles at lower tiers has minimal impact on inference cost when lower-tier models are at least 50×50$\times$50 × cheaper than the largest model. In such cases, the relative cost $(\gamma)$ satisfies $\gamma\leq\frac{1}{50}$, making ABC effective without parallelization. Finally, §5.2 explores practical scenarios where low relative costs make ABC is naturally suited for real-world deployments.
@@ 5.1.1 When Parallelization is Cheap
Figure 3: Fraction of inference cost saved as a function of relative cost of models $(\gamma)$, assuming a fixed selection rate $\mathbb{P}(r(x)=0)$. As parallelization decreases from fully parallel $(\rho=1)$ to sequential $(\rho=0)$, cost of evaluating ensembles increase and the cost savings decrease. When models across tiers are of similar size (e.g., smaller model is at most 5×5$\times$5 × smaller, $\gamma\geq\frac{1}{5}$), some parallelization is needed for ABC to reduce costs effectively. However, for lower relative costs (e.g., smaller model is at least 50×50$\times$50 × smaller, $\gamma\leq\frac{1}{50}$), sequential and parallel settings achieve meaningful savings, showing ABC ’s efficiency even with the added cost of using ensembles.
The accuracy vs. FLOPs for ABC is shown in Figure 2. As a point of comparison, we also include Wisdom-of-Committees (WoC) (Wang et al., 2021), a popular and representative confidence-based model cascading method. In most cases, we observe that ABC is able to improve the Pareto frontiers, as it usually sees a 111222 point increase in accuracy. We attribute this to a combination of (a) the improvement ensembles can have on accuracy that is widely noted in literature (Gontijo-Lopes et al., 2022; Jiang et al., 2023) and (2) the improvement agreement-based rules can cause in cascading (Appendix A). Overall, in terms of accuracy, ABC either exceeds—or at least matches—the accuracy of the best models at a fixed FLOPs budget, and is a practical drop-in replacement for these models.
@@ 5.2 Real-world Use-cases
[…]
@@ 5.2.1 Communication Cost in Edge-to-Cloud Inference
Our results, as shown in Figure 4, show that the flexibility that ABC affords in terms of model placement allows significant latency reductions, while providing superior accuracy compared to a single cloud model. In particular, we see that cascading in these scenarios provides an  14$\times$ reduction in communication cost for language tasks like SST-2; and for image datasets, we see a 5×5$\times$5 × reduction for ImageNet-1K and a 8×8$\times$8 × reduction in CIFAR10.
Takeaway #2: ABC enables model placement flexibility where small models run locally and large models in the cloud. Communication delays can make relative costs significant, resulting in ABC achieving 5-14$\times$ reductions in communication costs while maintaining superior accuracy.
@@ 5.2.2 Monetary Cost for Model Serving on Heterogeneous Hardware
Takeaway #3: GPU rental cost differences significantly exceed throughput differences. ABC optimally places models across hardware tiers, achieving 3$\times$ cost reductions for image tasks and 10-30% savings for text tasks.
@@ 5.2.3 Monetary Cost for Black-box API-Based Inference
Finally, in the current landscape dominated by LLMs, many providers offer black-box API access to their proprietary LLMs (Abdalla et al., 2023; Sun et al., 2022; Hadi et al., 2024). Similarly to the large cost difference between GPU generations discussed in the previous subsection, we also observe a large cost disparity between various generations/tiers of API calls. For example, using Together.ai—one of the lowest-cost serverless endpoint providers for LLMs111together.ai/pricing as of September 2024.—as a case study, we find that models within the 7B-8B range cost \$0.20 per million tokens, while LlaMA3.1-405B costs \$5.00. This translates to the larger model costing 25$\times$ more than the smaller model range $(\gamma=\frac{1}{25},\rho=1)$. If we consider GPT-4 as the gold standard, the cost of usage quickly scales to 150$\times$ of our reference smaller model’s costs. [^2] GPT-4-1106-preview costs \$30 as of September 2024..
@@ 5.3 Ablation Studies and Sensitivity Analysis
§5.1.2 and Figure 3 examine how ABC depends on the relative cost $\gamma$ between cascade tiers. When models have similar costs ($\gamma\geq\frac{1}{5}$), some parallelization may be required for cost savings. However, when smaller models are at least 50$\times$ cheaper ($\gamma\leq\frac{1}{50}$), ABC provides meaningful savings even with sequential execution, validating our theoretical analysis.
@@
Appendix B and Figure 6 demonstrate that agreement threshold $\theta$ estimation is stable across different model accuracy levels (37.6% to 86.0%) and converges with minimal validation data. Using only 100 samples yields threshold estimates that remain stable when evaluated on 10$\times$ more data, confirming the effectiveness of the calibration procedure.
@@ Appendix B Estimating Agreement Threshold $\theta$
Figure 6: Estimation of agreement threshold $\theta$ stability as a function of the number of samples used, across different model accuracy levels on the ImageNet-1K dataset. Each plot corresponds to a model with a specific accuracy, shown in the legend. The initial estimate is with 100100100 samples, with subsequent estimates with larger and larger number of samples. The initial estimate is reasonably close within subsequent estimates with larger number of samples.
@@ D.2 Details for Black-Box API Experiments
  • AutoMix: AutoMix trains a different router for all possible cascade steps, i.e., $n-1$ routers for the n cascade tiers, and this has to be repeated for every new task or model replacement in the system. A threshold (AutoMix+T) or POMDP (AutoMix+P) is trained with a combination of 50$\geq$ 50≥ 50 training samples from the same test data distribution and the initial inferences generated on the test data by the two models involved in each cascading step. After training the router, using the cascading system often involves running a few-shot self-verification 8 times at a high sampling temperature (temp = 1.0), using the same model that generates inference at the given cascade tier. Automix then averages the self-verification results and meta-verify it with the best parameters of the routing strategy to decide when to exit.
  • FrugalGPT: Just like AutoMix, FrugalGPT needs to train $n-1$ routers for $n$ cascade tiers, and each router needs to have a sense of the data distribution and the model’s predictive power. 500$\geq$ 500≥ 500 training samples and inference generated on these samples by the tier’s model are needed to train each tier/model’s scorer, a DistilBERT (Sanh et al., 2019).
@@
As shown in Figure 5, we observe that ABC is often more cost-effective than the baselines, even with their sophisticated routing mechanisms and singular model tiers. Our analyses show that ABC’s efficient deferral strategy allows a more aggressive utilization of cheaper models for a significant portion of the input while only using expensive models when necessary. For instance, we realize—upon analyzing FrugalGPT’s results—that the trained scorer struggles as an efficient deferral signal as the tasks get harder; hence, it is more likely to take the safer option to cascade as test sample difficulty increases. This means that ABC can be expected to be more efficient since the scaling laws ensure that the sum of the costs of using several cheaper models is still much less than the cost of using the larger model in the next tier. AutoMix, on the other hand, uses a few-shot self-verification system that is sampled >1>1> 1 times; hence, the additional API calls add significantly to its cost of usage. Considering that the self-verification process is an integral part of the AutoMix setup, it can be guaranteed that ABC will always be cheaper to use than AutoMix, despite using more models.
@@ E.2 Cost Benefits
Table 4: GPU rental costs from Lambda Cloud (Lambda, 2024) (September 2024) showing the substantial cost disparities between hardware generations. The 25$\times$ cost difference between H100 and V100 GPUs, combined with more modest throughput differences, creates favorable conditions for ABC’s heterogeneous hardware placement strategy described in Section 5.2.2

### v4

@@ Abstract
Adaptive inference schemes reduce the cost of machine learning inference by assigning smaller models to easier examples, attempting to avoid invocation of larger models when possible. In this work we explore a simple, effective adaptive inference technique we term Agreement-Based Cascading (ABC). ABC builds a cascade of models of increasing size/complexity and uses agreement between ensembles of models at each level of the cascade as a basis for data-dependent routing. Although ensemble execution introduces additional expense, we show that these costs can be easily offset in practice due to large expected differences in model sizes, parallel inference execution capabilities, and accuracy benefits of ensembling. We examine ABC theoretically and empirically in terms of these parameters, showing that the approach can reliably act as a drop-in replacement for existing models and surpass the best single model it aims to replace in terms of both efficiency and accuracy. Additionally, we explore the performance of ABC relative to existing cascading methods in three common scenarios: (1) edge-to-cloud inference, where ABC reduces communication costs by up to $14\times$; (2) cloud-based model serving, where it achieves a $3\times$ reduction in rental costs; and (3) inference via model API services, where ABC achieves a $2$-$25\times$ reduction in average price per token/request relative to state-of-the-art LLM cascades.
@@ 1 Introduction
  • We empirically evaluate ABC on a wide range of image and language tasks and find that ABC not only improves efficiency, but also accuracy, compared to the model that it aims to replace. We then consider the performance of ABC relative to existing cascading methods in common inference scenarios, including (1) edge-to-cloud inference where ABC reduces communication costs by up to $14\times$, (2) model-serving on heterogeneous GPUs, where ABC reduces rental costs by up to $3\times$ and (3) inference using black-box access to model API services, where ABC shows up to a $25\times$ reduction in average price per token.
@@ 4.2 Deferral Rule-Based Drop-in Cascade
A deferral-based cascade consists of a finite set of models and a deferral rule. Here, the idea is to start with the most resource-efficient model and use the deferral rule to determine if a better-performing model with a higher resource cost should be used for the current sample. The deferral rules themselves are designed to have negligible inference cost given the models inference outputs. For this exposition, we will restrict ourselves to cascades with only two levels, though the discussion in this section can be readily generalized to larger cascades. Let the cascade consist of an ensemble $H^{k}_{1}$ at the lower level, and the larger model $h_{2}$ at the higher level; $\mathcal{M}={\{H^{k}_{1},h_{2}\}}$. Let $r(x)$ denote a deferral rule;
@@
In this work, we focus on deployment scenarios where a decrease in model performance is a critical concern. This is particularly relevant when cascades are intended to function as a drop-in replacement for an existing deployment of an expensive model. This notion can be formulated as maximizing the number of calls to the smaller model while retaining the accuracy of the larger model. With a slight abuse of notation, let $\mathcal{M}_{r}(x)$ denote the prediction of the cascade with $\mathcal{M}={\{H^{k}_{1},h_{2}\}}$ and deferral rule $r$. Then we desire that for a choice of a small error budget $\xi>0$,
@@ 4.3 Deferral using Ensemble Agreement
Assuming access to a safe deferral rule, observe that for $\xi\geq\epsilon>0$, such rules are feasible for the program in Equation 2 (see Proposition 4.1). This implies that every cascade, $\mathcal{M}={\{H_{1}^{k},h_{2}\}}$ using a safe deferral rule $r$, is competitive with the larger model $h_{2}$ in terms of accuracy. We empirically evaluate these drop-in cascades in §5. Note that the optimal safe deferral rule can lead to an improvement in the accuracy of ABC in theory (see, Appendix A), and we see accuracy improvements in our experiments (§5).
@@
$r_{\text{v}}(x;\theta_{v})$ $=\begin{cases}1&\text{vote}(x;H^{k}_{1})\leq\theta_{v}\\ 0&\text{otherwise}.\end{cases}$ (3)
$r_{\text{s}}(x;\theta_{s})$ $=\begin{cases}1&s(x;H^{k}_{1})\leq\theta_{s}\\ 0&\text{otherwise}.\end{cases}$ (4)
@@ 4.4 Inference Cost Savings and Competitiveness
Let $\mathcal{M}={\{H_{1}^{k},h_{2}\}}$ be two classifiers and $r$ a deferral rule such that $r$ is a safe deferral rule for $H_{1}^{k}$ according to Definition 4.1, for a distribution $\mathbb{P}$ over $\mathcal{X}\times\mathcal{Y}$. Then for every $\xi\geq\epsilon>0$, the agreement based cascading (ABC) classifier $\mathcal{M}_{r}(x)$ is such that,
@@
The proof follows directly from Definition 4.1 and basic probability (see Appendix A). Safe deferral rules can be constructed with minimal cost: our threshold estimation requires only $\sim$100 validation samples and simple voting computations, avoiding expensive router training. Thus, the cost savings we can expect with a drop-in cascade depend on three key factors: the relative cost $\gamma$, the degree of parallelization $\rho$ and the deferral rate $P(r(x)=1)$, or equivalently, the selection rate $P(r(x)=0)$. In particular, for every feasible deferral rule $r$, in the best case scenario where the cost of the smaller model is negligible (i.e., $\gamma=0$), the cost of inference reduces by the selection rate, $P(r(x)=0)$. Conversely, in the worst-case scenarios, the cost can be $(k+1)$ times the cost of the larger model. In the next section, we demonstrate real world scenarios where the favorable interplay of these three quantities lead to significant improvements in inference cost. See Appendix C and Figure 7 for information on selection rates for various models and datasets considered here.
@@ 5 Experiments
This section evaluates our training-free Agreement-Based Cascading (ABC) approach across a variety of language and vision tasks, focusing on accuracy and inference cost. First, in §5.1.1, we examine ABC’s accuracy-cost tradeoff against state-of-the-art models under full parallelization, setting the parallelization factor $(\rho)$ to $\rho=1.0$ for optimal tradeoff. Next, in §5.1.2, we show that using ensembles at lower tiers has minimal impact on inference cost when lower-tier models are at least $50\times$ cheaper than the largest model. In such cases, the relative cost $(\gamma)$ satisfies $\gamma\leq\frac{1}{50}$, making ABC effective without parallelization. Finally, §5.2 explores practical scenarios where low relative costs make ABC is naturally suited for real-world deployments.
@@ 5.1.1 When Parallelization is Cheap
Figure 3: Fraction of inference cost saved as a function of relative cost of models $(\gamma)$, assuming a fixed selection rate $\mathbb{P}(r(x)=0)$. As parallelization decreases from fully parallel $(\rho=1)$ to sequential $(\rho=0)$, cost of evaluating ensembles increase and the cost savings decrease. When models across tiers are of similar size (e.g., smaller model is at most $5\times$ smaller, $\gamma\geq\frac{1}{5}$), some parallelization is needed for ABC to reduce costs effectively. However, for lower relative costs (e.g., smaller model is at least $50\times$ smaller, $\gamma\leq\frac{1}{50}$), sequential and parallel settings achieve meaningful savings, showing ABC ’s efficiency even with the added cost of using ensembles.
The accuracy vs. FLOPs for ABC is shown in Figure 2. As a point of comparison, we also include Wisdom-of-Committees (WoC) (Wang et al., 2021), a popular and representative confidence-based model cascading method. In most cases, we observe that ABC is able to improve the Pareto frontiers, as it usually sees a 1122 point increase in accuracy. We attribute this to a combination of (a) the improvement ensembles can have on accuracy that is widely noted in literature (Gontijo-Lopes et al., 2022; Jiang et al., 2023) and (2) the improvement agreement-based rules can cause in cascading (Appendix A). Overall, in terms of accuracy, ABC either exceeds—or at least matches—the accuracy of the best models at a fixed FLOPs budget, and is a practical drop-in replacement for these models.
@@ 5.2 Real-world Use-cases
As noted in the previous two subsections, ABC  either improves on, or is at least competitive with, the single best model in terms of accuracy. Moreover, whenever the relative cost is small enough, ABC  suffers negligible additional penalty for using an ensemble of models. As noted in Figure 1, accuracy vs inference cost scaling for ML models already imply that the relative cost of models that only differ a few points in accuracy is small. For instance, the state-of-the-art model performance on the ImageNet-1K dataset attains about $83\%$ top-1 accuracy with 7070B FLOPs. A Pareto-optimal model that achieves $63\%$ accuracy requires only about 11B flops, and thus a two-level ABC with ensembles of these two models has a relative cost of $\gamma=\frac{1}{70}$.
@@ 5.2.1 Communication Cost in Edge-to-Cloud Inference
[…]
@@ 5.2.2 Monetary Cost for Model Serving on Heterogeneous Hardware
[…]
@@ 5.2.3 Monetary Cost for Black-box API-Based Inference
[…]
@@ 5.3 Ablation Studies and Sensitivity Analysis
[…]
@@
[…]
@@ 6 Conclusion
Acknowledgements
We thank Xiaofang Wang, Wittawat Jitkrittum, Lucio Dery, Pratiksha Thaker, and Kevin Kuo for their helpful comments. This work was supported in part by the National Science Foundation grants IIS2145670, CCF2107024, IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Amazon, Apple, Google, Intel, Meta, Morgan Stanley, Scribe, and the CyLab Security and Privacy Institute. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
@@ Appendix B Estimating Agreement Threshold $\theta$
Figure 6: Estimation of agreement threshold $\theta$ stability as a function of the number of samples used, across different model accuracy levels on the ImageNet-1K dataset. Each plot corresponds to a model with a specific accuracy, shown in the legend. The initial estimate is with $100$ samples, with subsequent estimates with larger and larger number of samples. The initial estimate is reasonably close within subsequent estimates with larger number of samples.
@@ D.2 Details for Black-Box API Experiments
  • AutoMix: AutoMix trains a different router for all possible cascade steps, i.e., n1n-1 routers for the n cascade tiers, and this has to be repeated for every new task or model replacement in the system. A threshold (AutoMix+T) or POMDP (AutoMix+P) is trained with a combination of $\geq 50$ training samples from the same test data distribution and the initial inferences generated on the test data by the two models involved in each cascading step. After training the router, using the cascading system often involves running a few-shot self-verification 8 times at a high sampling temperature (temp = 1.0), using the same model that generates inference at the given cascade tier. Automix then averages the self-verification results and meta-verify it with the best parameters of the routing strategy to decide when to exit.
  • FrugalGPT: Just like AutoMix, FrugalGPT needs to train n1n-1 routers for $n$ cascade tiers, and each router needs to have a sense of the data distribution and the model’s predictive power. $\geq 500$ training samples and inference generated on these samples by the tier’s model are needed to train each tier/model’s scorer, a DistilBERT (Sanh et al., 2019).
@@
As shown in Figure 5, we observe that ABC is often more cost-effective than the baselines, even with their sophisticated routing mechanisms and singular model tiers. Our analyses show that ABC’s efficient deferral strategy allows a more aggressive utilization of cheaper models for a significant portion of the input while only using expensive models when necessary. For instance, we realize—upon analyzing FrugalGPT’s results—that the trained scorer struggles as an efficient deferral signal as the tasks get harder; hence, it is more likely to take the safer option to cascade as test sample difficulty increases. This means that ABC can be expected to be more efficient since the scaling laws ensure that the sum of the costs of using several cheaper models is still much less than the cost of using the larger model in the next tier. AutoMix, on the other hand, uses a few-shot self-verification system that is sampled >1>1 times; hence, the additional API calls add significantly to its cost of usage. Considering that the self-verification process is an integral part of the AutoMix setup, it can be guaranteed that ABC will always be cheaper to use than AutoMix, despite using more models.
@@ E.2 Cost Benefits
[…]